Find the eigenvalues and associated eigenfunction of the strum-Liouville problem xy"+y'+lamda/2 y=0,y'(1)=0,y'(e^2π)=0
1
Expert's answer
2020-08-16T20:58:27-0400
"Solution"

"xy''+y'+ \frac{\lambda}{2 }y=0"



"y'(1)=0,y'(e^2\u03c0)=0"


We begin by making some assumptions which will simplify the problem. This will turn our differential equation into a constant-coefficient equation.

Let

"t=lnx"

And

"y(x)=\phi (ln x)=\phi(t)"


"y'(x)=\frac{1}{x}(\frac{d \phi (t)}{dt})"


"y''(x)=\frac{1}{x^2}(\frac{d^2 \phi (t)}{dt^2}-\frac{d \phi (t)}{dt})"


Next, we substitute our new y's to the original equation and simplify until we get a second order, constant coefficient equation.


"x(\frac{1}{x^2}(\frac{d^2 \phi (t)}{dt^2}-\frac{d \phi (t)}{dt}))+\frac{1}{x}(\frac{d \phi (t)}{dt})+\frac{\lambda}{2 }\phi(t)=0"


"(\frac{d^2 \phi (t)}{xdt^2}-\frac{d \phi (t)}{xdt})+(\frac{d \phi (t)}{xdt})+\frac{\lambda}{2 }\phi(t)=0"


"\frac{d^2 \phi (t)}{xdt^2}+\frac{\lambda}{2 }\phi(t)=0"

This is easy to solve with a characteristic equation.


"0=r^2+\frac{\lambda}{2}x"

Solving for r we get

"\sqrt{r^2}=\sqrt{-\frac{\lambda}{2}x} \implies r=\sqrt{-\frac{\lambda}{2}x}"

We now write out the solution in exponential form.


"\phi(t)=C_1 \epsilon^{(\sqrt{\frac{-\lambda x}{2}})t}+C_2 \epsilon^{(\sqrt{\frac{-\lambda x}{2}})t}"

Converting back to "y(x)" we get


"y(x)=C_1 \epsilon^{(\sqrt{\frac{-\lambda x}{2}})ln x}+C_2 \epsilon^{(\sqrt{\frac{-\lambda x}{2}})ln x}"

Applying Euler's identities and picking two linear independent solutions


"y(x)=C_3 \epsilon^{0}sin(\sqrt{\frac{-\lambda x}{2}})ln x+C_4 \epsilon^{0} cos (\sqrt{\frac{-\lambda x}{2}})ln x"

"\epsilon ^{0} = 1"


"y(x)=C_3 sin(\sqrt{\frac{-\lambda x}{2}})ln x+C_4 cos (\sqrt{\frac{-\lambda x}{2}})ln x"

Substituting our boundary conditions.


"y=0,y'(1)=0,y'(e^2\u03c0)=0"

Starting with y'(1) we get


"y(1)=C_3(1)sin(\sqrt{\frac{-\lambda x}{2}ln(1)}+C_4(1)cos(\sqrt{\frac{-\lambda x}{2}ln(1)}"

"y(1)=0=C_3sin0+C_4cos0=C_4 \implies C_4=0"


"y(x)=C_3(1)sin(\sqrt{\frac{-\lambda x}{2}ln(1)}"

"y'(\epsilon ^{2\pi})=0"


"y'(\epsilon^{2\pi})=0=C_3(\epsilon^{2\pi})sin(\sqrt{\frac{-\lambda x}{2}ln(\epsilon^{2\pi})}"


"n \pi=\sqrt{\frac{-\lambda x}{2}ln(\epsilon^{2\pi})} \implies \lambda=(\frac{2n\pi}{x \epsilon^{2\pi}})^2--->Eigenvalue"


Simplifying to get the eigenfunction, we get


"y(x)=Cx\sin \sqrt{x(\frac{2n\pi}{2x \epsilon^{2\pi}})^2}lnx \implies y(x)=Cx\sin \frac{n\pi}{x \epsilon^{2\pi}}\sqrt{lnx}--->Eigenfunction"